OpenStudy (anonymous):

Anyone know a simple way to remember how to get slopes? Anyone know a simple way to remember how to get slopes? @Mathematics

OpenStudy (anonymous):

rise over run

OpenStudy (anonymous):

find dy/dx

OpenStudy (anonymous):

I have a great way, The steeper the hill, the harder it is for you to keep yourself moving Keeping this fact in mind, by definition, the slope is the measure of the steepness of a line. In real life, we see slope in any direction. However, in math, slope is defined from left to right. I repeat we always measure slope going from left to right. This is very important! There are three types of slope you can encounter. A slope can be positive, negative, or equal to zero. When the slope is equal to zero, we say that there is no slope A positive slope: If you go from left to right and you go up, it it a positive slope

OpenStudy (anonymous):

A negative slope: If you go from left to right and you go down, it it a negative slope

OpenStudy (anonymous):

A zero slope: If you go from left to right and you don't go up or down, it it a zero slope

OpenStudy (anonymous):

How is that?

OpenStudy (anonymous):

if u know any point of grafic, you may use: (y-y0)= m(x - x0) where y0 and x0 are the coordinates of the known point and m is the slope.

OpenStudy (anonymous):

y-y0 so slope= ------ x - x0

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome

myininaya (myininaya):

you could think about what speed equals speed=distance over time s=d/t and easy way to remember this ratio is recalling the units in which you drive well us americans say miles per hour miles is talking about the distance and per is division hours is talking about time you can think of if like this y is the change in distance and x is the change in time s=d/t=change of distance/change of time=change of y/change of x

myininaya (myininaya):

we do say speed is positive and we are looking for the rate of change (aka slope) so don't do the absolute value of

myininaya (myininaya):

well positive or neutral*